### v2

@@ 3.2 AGV scheduling: ILP model
Concerning order variables, for each pair of AGVs passing a zone there is a pair of these ($y_{in}$ and $y_{out}$). However, the order of AGVs cannot change in a zone (this will be enforced by constraints in Eq. (22)), a single order variable per zone and AGVs’ pair suffices; there are $|J|\left(|J|-1\right)/2$ such pairs. Further, if hypothetically the whole topology consisted of bidirectional lanes, we had #$\#z=\#y$. In a more general layout, inequality #$\#z\leq\#y$ holds. Hence,
@@
The minimal headway constraint (mh) is the model input that determines minimal time spacing between two subsequent AGVs. In detail, consider two AGVs $j,j^{\prime}$ heading in the same direction. Let $s*,s,s^{\prime}$ be the sequence of subsequent zones they both pass. Then:
@@
Here $M$ is a large enough number for inequality to always hold for $y(j^{\prime},j,s)=1$ (we used so called “big M encoding”). If $s$ does not have a successor in the sequence, we do not consider Eq. (15). Analogously , if $s$ does not have the predecessor, we do not consider Eq. (16).
Refer to captionRefer to caption
@@
In the scheduling theory language: the assumption that the zone can be occupied by one AGV only defines the job shop machine environment, and the prescription of the passing time over zones imposes processing time constraints.
@@
To approximate the number of zone constraints (zc), the upper limit is obtained again with the hypotatical assumption that all AGVs pass all zones. Then for each zone and each AGV we have $|J|-1$ inequalities from Eq. (19) and one inequality from Eq. (20), yielding
@@
is also expected to be linear in $|S|$ and quadratic in $|S|$.
Altogether the problem size (in terms of the number of constraints and binary variables) scales linearly in the number of zones $|S|$ and quadratically in the number of AGVs $|J|$. The quadratic scaling can cause difficulties when applying the ILP model to a large number of AGVs. As such, this AGV problem seems to be a good use-case for the quantum approach described at the beginning of this section.
@@ 4 Computational results and discussion
For numerical calculations we used examples of problems similar to those presented in Fig. 1, but with different numbers of AGVs and zones. This is a typical setting in an industrial system, as various AGVs are in traffic at various times, and collision zones form dynamically. As a typical example of the particular parameter setting with $7$ zones see Fig. 3. The actual values of parameters can be read from the topology, AGVs’ speeds, and zone locations. The example with $7$ zones and $7$ AGVs will be discussed in more detail in Appendix B.
Figure 3: Example network with $7$ zones derived from Fiq. 1
From an optimization point of view, one of our goals is to examine the scaling of the problems’ computational time and number of variables and constraints with the size of the problem in terms of $|J|$ and $|S|$ (namely the number of AGVs and zones). We start with $2$ AGVs and $4$ zones and end with $15$ AGVs and $7$ zones. The sizes of the instances we have solved are tabulated in Tab. 1. Referring to somehow analogous problem in railways in [36] with $3$ to $5,5$ constraints per variable the AGV problem is much more sparse (approximately $2$ constraints per variable). Such less dense problem is more probable to be better handled on the quantum device, where the embedding on the device with limited degre connectivity is performed.
@@
2 / 4 / 10 16  12/4 4 16 20 16 32
4 / 4 / 10 36  20/16 10 34 56 62 124
6 / 7 / 40 78  38/40 32 80 189 252 504
7 / 7 / 40 118  48/70 55 127 245 343 686
12 / 7 / 40 596  102/494 429 766 630 1008 2016
15 / 7 / 40 696  118/576 495 883 945 1575 3150
Table 1: Sizes of ILPs of the analyzed AGV problems. The notion of upper limits is described in Section 3.2. We have approximately $2$ constraints per variable (equality and inequality), what makes the the problem sparse in comparison with railway counter example in [36].
Figure 4: Comparison of performance of Classical CPLEX exact and approximate (achieved by setting constant computational time) with hybrid quantum-classical solver, CQM in particular. All presented solutions were feasible. As concerning BQM feasible solutions were found only for small instances.
All classical (that is not quantum) calculations were performed on the consumer-grade computer described in Tab. 2, and quantum calculations were performed on DWave’s Advantage_systems6.2 quantum annealer, whose physical properties are described in [34]. In short, the quantum device has over $5600$ qubits and $40$ thousand couplings between them, organized into the Pegasus topology of QPU. Concerning pure quantum computing, we were able to fit to real quantum device only first two smallest problems, from Table. 1. The larger of $36$ ILP variables and $44$ ILP constrains (equalities and inequalities) was converted into $268$ qubits Ising problem with $2644$ (quadratic) couplings between. However, even for these two small problems we did not achieve the feasible solution. From this, we can conclude that quantum annealers are too small and too prone to errors, and hence for more accurate solutions of larger problem we opt for hybrid quantum - classical solvers.
@@
Figure 5: Histograms of objectives of feasible CQM solutions for various problem sizes: $6$ AGVs a), $7$ AGVs b), $12$ AGVs c), $15$ AGVs d). For each run approx hundred of samples were returned. Histogram spreads are measured by standard deviation (std): $6$ AGVs std = $0.19$, $7$ AGVs std = $0.68$, $12$ AGVs std = $1.66$, $15$ AGVs std = $2.08$.
@@
Figure 6: Details of hybrid solvers. The QPU time is in the left panel, and the number of broken constraints of the solution is in the right panel. Observe that the output of the CQM solver was satisfactory, which was not the case in the BQM solver, especially for large problems.
Our first approach was to use the BQM solver. In order to do so, the ILP has to be converted to QUBO. Appendix A describes the details of this transcription. The constraints were taken into account with penalties using ad-hoc penalty coefficients. We have found that except for the smallest problems, the BQM solver does not return high-quality solutions (see Fig. 6, right panel). Notably, the solutions were not feasible: a more advanced and systematic determination should replace our ad-hoc choice of the penalty coefficients. As this issue is handled in the CQM solver, we switched to that one. Indeed, the CQM hybrid solver, which is convenient also because it accepts ILPs directly, has always provided some feasible solutions in our problems.
@@
We have therefore made a comparison of 3 approaches: the exact CPLEX ILP solver wich always solves the problem to optimality, an approximate CPLEX solution, achieved by setting constant computational time for CPLEX solver, and the CQM hybrid solver as an approximate heuristic. Our comparison is summarized in Fig. 4. In terms of computational time, CQM has outperformed the classical exact CPLEX solver for large problems ($596$ and $696$ linear variables). It has to be noted that in each of the CQM computations, a nonzero QPU time was reported (see Fig. 6, left panel), indicating that the QPU was active during the calculation. The largest problem was of $15$ AGVs and $7$ zones, which is the sound size concerning the real live factory environment. As for the quality of the solutions, slightly higher objective value than the exact solution. This overview shows that the CPLEX approximate solver, performing all computations in $1$ second, is fast enough for almost online scheduling and dynamics response of our algorithm. Hence, such results would be admissible for practical application in real-life factories. In the future, when better quantum annealers will be available, they may become more efficient than classical heuristics. Concerning railway counterpart problem in  [36] and analogous real live problems of $817$ variables and over $3000$ constraints the performance of the CQM solver therein could compete only with exact CPLEX in terms of computational time but at larger cost of the solution quality than in the AGV case.
@@
As the CQM solver returns multiple feasible solutions at a time, in Fig. 5 we present the histograms of these, for problems of various sizes. Observe that for small problems, the results concentrate around the optimum, while for large problems, the histogram spreads towards higher-than-optimum objective values. It is important to note that even for large problems there are multiple feasible and potentially valuable solutions obtained. The results with the CQM solver are competitive and illustrate an efficient application of hybrid algorithms. The details of a particular case and its solutions, including a detailed comparison of classical and different quantum solutions are presented in Appendix B.
@@
For the largest problem of $15$ AGVs and $7$ zones ($696$ linear variables), the problem is still tractable using classical heuristics. As we expect the size of the problem to scale quadratically with the number of AGVs, other heuristics may be required to handle larger problems in expanding industrial environments; a hybrid quantum-classical approach may be one of these. This however would also require the improvement of the hybrid solvers as the feasibility percentage of CQM decreases with the problem size, as it can be observed in Fig. 6, middle panel. If the linear scaling between feasibility percentage and the number of linear variables holds, the biggest problems that can be solvable using the current CQM can have approximately up to $1000$ variables.

### v3

@@ 3.2 AGVs scheduling: ILP model
Concerning order variables, for each pair of AGVs passing a zone there is a pair of these ($y_{in}$ and $y_{out}$). However, the order of AGVs cannot change in a zone (constraints in Eq. (22) will enforce this), a single order variable per zone and AGVs’ pair suffices; there are $|J|\left(|J|-1\right)/2$ such pairs. Further, if hypothetically the whole topology consisted of bidirectional lanes, we had #$\#z=\#y$. In a more general layout, inequality #$\#z\leq\#y$ holds. Hence,
@@
The minimal headway constraint (mh) is the model input that determines minimal time spacing between two subsequent AGVs. In detail, consider two AGVs $j,j^{\prime}$ heading in the same direction. Let $s^{*},s,s^{\prime}$ be the sequence of subsequent zones they both pass. Then:
@@
Here, $M$ is a large enough number for inequality to always hold for $y(j^{\prime},j,s)=1$ (we used the so-called "big M encoding"). If $s$ does not have a successor in the sequence, we do not consider Eq. (15). Analogously, if $s$ does not have the predecessor, we do not consider Eq. (16).
@@
In the scheduling theory language: the assumption that one AGV can occupy the zone only defines the job shop machine environment, and the prescription of the passing time over zones imposes processing time constraints.
@@
To approximate the number of zone constraints (zc), the upper limit is obtained again with the hypothetical assumption that all AGVs pass all zones. Then for each zone and each AGV we have $|J|-1$ inequalities from Eq. (19) and one inequality from Eq. (20), yielding
@@
is also expected to be linear in $|S|$ and quadratic in $|J|$.
Altogether, the problem size (in terms of the number of constraints and binary variables) scales linearly in the number of zones $|S|$ and quadratically in the number of AGVs $|J|$. The quadratic scaling can cause difficulties when applying the ILP model to a large number of AGVs. As such, this AGV problem seems to be a good use-case for the quantum approach described at the beginning of this section.
3.3 QUBO model
The most general form of inequalities in our model is in Eq. (15) or Eq. (16). Based on these, we present a detailed path of transformation of these inequalities into QUBO. Let us contract $t$-variables and $y$-variables to vectors with elements $t_{j}$ and $y_{i}$ (and define the respective vector for the constant terms $c_{i}$). Then, from the aforementioned equations , we have:
$t_{j}-t_{j^{\prime}}-My_{i}+\xi_{i}=-c_{i}.$ (29)
To replace inequalities with equalities, we use slack variables
as $\min(t_{j})=v_{j}$, and $\max(t_{j})=v_{j}+d_{\max}$ from Eq. (5), and $\upsilon_{j^{\prime}}=\upsilon_{j}+c_{i}$. Observe that some of the inequalities such as these of the minimal passing time constraints in Eq. (14) do not have order variables ($y$ or $z$), for such variables $\bar{\xi}_{i}=d_{\max}$.
To convert the constrained problem in Eq. (29) into unconstrained, we use the penalty method: each constraint violation is penalized by the hard constraint penalty $p>0$. Such a penalty has to be sufficiently large not to be overruled by the objective. Then, the optimization problem has the form:
and referring to Eqs. (10) (12) (27):
#$\#\text{QUBO}\leq 2|J||S|\big{\lceil}\ln_{2}(d_{\text{max}}+1)\big{\rceil}+3|J|^{2}|S|\big{\lceil}\ln_{2}(d_{\text{max}}+\max_{i}\xi_{i}+1)\big{\rceil}+\frac{|J|\left(|J|-1\right)|S|}{2}=O(|J|^{2}|S|).$ (33)
The size of the problem (in terms of the number of QUBO variables) scales linearly with the number of zones and quadratically in the number of AGVs. The number of constraints per variable (the mean vertex degree of the graph model) is tied to the expansion of quadratic terms in Eq. (31). In Table. 1, the sizes of actual problems in terms of Ising variables (derived directly from QUBO) are presented. Hence, this is the number of the so-called logical qubits. Upon the procedure of using a hardware annealer, the problem has to be embedded into the physical topology of the annealer, which is a nontrivial procedure and is not always possible [44]. The actual problem solved on the hardware is thus expressed in terms of physical qubits, and a logical qubit can possibly be implemented using multiple physical qubits.
Apart from the two smallest problems from Table. 1, the others did not embed into the Pegasus chip, see Fig. 6. For larger instances, the hybrid, quantum-classical approach is necessary. For example, for the $15$ AGVs and $7$ zones problem we have $696$ variables (upper limit according to Eq. (12) was $945$), and $1378$ linear constraints (upper limit according to Eq. (28) (27) was $5513$).
$|J|$ / $|S|$
n.o. qubits
n.o. quadratic couplings
edges
edge density [45] $\frac{edges}{fullgraph}$ n.o. linear fields
2 / 4 122 1066 0.14 121
4 / 4 268 2644 0.07 267
6 / 7 796 11954 0.04 782
7 / 7 1204 19084 0.03 1183
12 / 7 6357 116422 0.006 6250
15 / 7 7343 134415 0.005 7219
21 / 7 13892 259702 0.005 13648
Table 1: Sizes of analyzed AGVs problems, in terms of Ising variables. Note, that the edge density is smaller than the one for a corresponding full graph, and it decreases with increasing problem size (up to 3 orders of magnitude for the largest considered problem).
@@ 4 Computational results and discussion
For numerical calculations, we used examples of problems similar to those presented in Fig. 1, but with different numbers of AGVs and zones. This is a typical setting in an industrial system, as various AGVs are in traffic at various times, and collision zones form dynamically. As a typical example of the particular parameter setting with $7$ zones see Fig. 3. The actual values of parameters can be read from the topology, AGVs’ speeds, and zone locations. The example with $7$ zones and $7$ AGVs will be discussed in more detail in Appendix A.
Figure 3: Example network with $7$ zones derived from Fig. 1
From an optimization point of view, one of our goals is to examine the scaling of the problems’ computational time, number of variables and constraints with the size of the problem in terms of $|J|$ and $|S|$ (namely the number of AGVs and zones). We start with $2$ AGVs and $4$ zones and end with $21$ AGVs and $7$ zones. The sizes of the instances we have solved are tabulated in Table. 2. It is interesting to compare this to an analogous problem in the railways solved by us in a previous contribution [39]: while there we had $3$ to $5.5$ constraints per variable, the AGVs problem is significantly sparser (approximately $2$ constraints per variable), making it more likely to be better handled by a quantum device, where the embedding on the device with limited degree connectivity is performed. Note, however, that if the considered AGVs problem address a more branched topology (e.g. with loops, etc.), this would result in a denser problem, more similar to the analogous railway-related one, with more constraints per variable. This would be less suited for the quantum-based approach.
@@
2 / 4 / 10 16  12/4 4 16 20 8 48
4 / 4 / 10 36  20/16 10 34 56 32 192
6 / 7 / 40 78  38/40 32 80 189 126 756
7 / 7 / 40 118  48/70 55 127 245 172 1029
12 / 7 / 40 596  102/494 429 766 630 504 3024
15 / 7 / 40 696  118/576 495 883 945 788 4725
21 / 7 / 40 1302  168/1134 1007 1705 1764 1544 9261
Table 2: Sizes of ILPs of the analyzed AGVs problems. The notion of upper limits is described in Section 3.2. We have approximately $2$ constraints per variable (equality and inequality), which makes the problem sparse in comparison with the railway counter-example in [39].
Figure 4: Comparison of performance of Classical CPLEX exact and approximate (achieved by setting constant computational time) with hybrid quantum-classical solver, CQM in particular. All presented solutions were feasible. For CPLEX exact on the largest instance ($1302$ variables) the time limit of $30$ minutes was used, as it was not possible to achieve the certified output in a reasonable time, $6$h time limit CPLEX computation yields the same results. Error bars were computed by performing $10$ independent realizations of experiments of CQM computation and calculating the standard deviation over realizations.
All classical (that is, not quantum) calculations were performed on the consumer-grade computer described in Table. 3, and quantum calculations were performed on D-Wave’s Advantage_systems6.2 quantum annealer, whose physical properties are described in [37]. In short, the quantum device has over $5600$ qubits and $40$ a thousand couplings between them, organized into the Pegasus topology of QPU. Concerning pure quantum computing, we were able to fit to real quantum device only the first two smallest problems from Table. 2. The particular sizes of these problems in terms of Ising variables are presented in Table. 1. The problem of $36$ ILP variables and $44$ ILP constraints (equalities and inequalities) was converted into $268$ qubits Ising problem with $2644$ (quadratic) couplings between. However, even for these two small problems, we did not achieve the feasible solution. From this, we can conclude that quantum annealers are too small and too prone to errors, and hence, for more accurate solutions to larger problems, we opt for hybrid quantum-classical solvers.
@@
Figure 5: Histograms of objectives of feasible CQM solutions for various problem sizes: $6$ AGVs a), $7$ AGVs b), $12$ AGVs c), $15$ AGVs d). For each run, approximately a hundred samples were returned. Histogram spreads are measured by the standard deviation (std): $6$ AGVs std = $0.19$, $7$ AGVs std = $0.62$, $12$ AGVs std = $1.51$, $15$ AGVs std = $1.9$.
@@
Figure 6: Details of hybrid solvers. The QPU time is in the left panel, the feasibility percentage is in the middle panel, and the number of broken constraints of the solution is in the right panel (pure QA - pure quantum annealing, SBM - simulated bifurcation). Observe that the output of the CQM solver was satisfactory in terms of feasibility for all instances.
@@
Indeed, the CQM hybrid solver, which is convenient also because it accepts ILPs directly, has always provided some feasible solutions in our problems. We have therefore made a comparison of 3 approaches: the exact CPLEX ILP solver which always solves the problem to optimality, an approximate CPLEX solution, achieved by setting constant computational time for CPLEX solver, and the CQM hybrid solver as an approximate heuristic. Our comparison is summarized in Fig. 4. In terms of computational time, CQM has outperformed the classical exact CPLEX solver for large problems ($596$ and $696$ linear variables). It has to be noted that in each of the CQM computations, a nonzero QPU time was reported (see Fig. 6, left panel), indicating that the QPU was active during the calculation. From the reported QPU computational times we can conclude that rather a small amount of computation was performed on the QPU, and the QPU time does not scale with the problem size. The probable explanation of this observation is that the D-Wave hybrid solvers use "Quantum Backstops". When stuck in a local minimum, the solver uses a quantum approach on selected sub-problem to move away from such a local minimum. The number of such "Quantum Backstops", and tied to them the QPU time, may not depend on the problem size. The largest problem was of $21$ AGVs and $7$ zones, which is the sound size concerning the real live factory environment. As for the quality of the solutions, the CQM yields slightly higher objective value than the exact solution. This overview shows that the CPLEX approximate solver, performing all computations in $1$ second, is fast enough for almost online scheduling and dynamics response of our algorithm. Hence, such results would be admissible for practical application in real-life factories. In the future, when better quantum annealers will be available, they may become more efficient than classical heuristics. Concerning railway counterpart problem in  [39] and analogous real live problems of $817$ variables and over $3000$ constraints, the performance of the CQM solver therein could compete only with exact CPLEX in terms of computational time but at larger cost of the solution quality than in the AGVs case.
@@
As the CQM solver returns multiple feasible solutions at a time, in Fig. 5, we present the histograms of these for problems of various sizes. Observe that for small problems, the results concentrate around the optimum. In contrast, for large problems, the histogram spreads towards higher-than-optimum objective values, the standard deviation value presented in caption grows by the order of magnitude, if we compare the $6$ AGVs problem with the $15$ AGVs one. It is important to note that even for large problems, there are multiple feasible and potentially valuable solutions obtained. The results with the CQM solver are competitive and illustrate an efficient application of hybrid algorithms. The details of a particular case and its solutions, including a detailed comparison of classical and different quantum solutions, are presented in Appendix A.
@@
For the largest problem of $21$ AGVs and $7$ zones ($1302$ linear variables), the problem is still tractable using classical heuristics. As we expect the size of the problem to scale quadratically with the number of AGVs, other heuristics may be required to handle larger problems in expanding industrial environments; a hybrid quantum-classical approach may be one of these. This, however, would also require the improvement of the hybrid solvers as the feasibility percentage of CQM decreases with the problem size, as it can be observed in Fig. 6, middle panel. If the linear scaling between feasibility percentage and the number of linear variables holds, the biggest problems that can be solvable using the current CQM can have approximately up to $1000$ variables.
Recall that our first approach was to use the BQM solver. In order to do so, the ILP has to be converted to QUBO. Section 3.3 describes the details of this transcription. The constraints were taken into account with penalties using ad-hoc penalty coefficients. We have found that except for the smallest problems, the BQM solver does not return high-quality solutions (see Fig. 6, right panel). Notably, the solutions were not feasible: a more advanced and systematic determination should replace our ad-hoc choice of the penalty coefficients. The interesting observation here is, however, that for BQM, the percentage of non-feasible solutions is constant for larger instances and equal to these from the pure Quantum Annealing for smaller ones. Hence, here the unfeasibility of Quantum Annealing at the lower scale may have an effect on the BQM performance on the larger scale. For the comprehensive analysis, in Fig. 6, the right panel we compare two QUBO base solvers, already mentioned proprietary, closed source BQM and open source Simulated Bifurcation Machine (SBM) solver. The SBM behaves better than the BQM, but still does not return feasible solutions for large instances. This observation paves the way for the development of new quantum-inspired solvers, as the algorithm behind SBM is known.
Drawing inspiration from quantum adiabatic optimization with a network of nonlinear quantum oscillators, SBM exploits the phenomenon of bifurcation in a system of classical, nonlinear system of Hamilton equations[46]:
$H$ $=\frac{a_{0}}{2}\sum_{i=1}^{N}p_{i}^{2}+\sum_{i=1}^{N}\left(\frac{q_{i}^{4}}{4}+\frac{a_{0}-a(t)}{2}q_{i}^{2}\right)-c_{0}\sum_{i=1}^{N}\left(h^{\prime}_{i}q_{i}+\frac{1}{2}\sum_{j=1}^{N}J^{\prime}_{ij}q_{i}q_{j}\right)$ (34)
$\dot{q}_{i}$ $=\frac{\partial H}{\partial p_{i}}=a_{0}p_{i}$ (35)
$\dot{p}_{i}$ $=-\frac{\partial H}{\partial q_{i}}=-\left[q_{i}^{2}+a_{0}-a(t)\right]q_{i}+c_{0}\left(\sum_{j=1}^{N}J^{\prime}_{ij}q_{j}+h^{\prime}_{i}\right)$ (36)
It can be interpreted as a motion of particles with mass $a_{0}^{-1}$, position $q_{i}\in\mathbb{R}$, and momentum $p_{i}\in\mathbb{R}$, in a time-dependent potential and coupled via Ising-like interaction. Matrix $J^{\prime}$ and vector $h^{\prime}$ encode the Ising minimization problem in question, cf. Eq(3). Numbers $a_{0}$ and $c_{0}$ are constant hyperparameters (in our calculations, we use $a_{0}=1$ and $c_{0}=1/\lambda_{\text{max}}$, where $\lambda_{\text{max}}$ is the largest eigenvalue of $J^{\prime}$), whereas $a(t)$ is a linearly increasing function of time, which drives the system through the bifurcation point, occurring approximately when $a(t)=a_{0}$. Post-bifurcation, the energy landscape governing the evolution of particles approximately encodes the local minima of the Ising term, and thus the particles flow towards the low-energy solutions of the optimization problem, which then can be extracted by taking the sign $s_{i}=\mathrm{sign}(q_{i})$. In an effort to suppress the errors originating from relaxing the variables from binary to continuous, a modified version of SBM was introduced[47] (which we use in this work), where the nonlinear term $q_{i}^{4}$ was replaced by perfectly inelastic walls, located at $|q_{i}|=1$, and the Ising contribution to dynamics was discretized, i.e. term $\sum_{j=1}^{N}J^{\prime}_{ij}q_{j}$ in Eq. (36) got replaced by $\sum_{j=1}^{N}J^{\prime}_{ij}\mathrm{sign}(q_{j})$. Additionally, these modifications guarantee that the dynamics can be stopped when $a(t)=a_{0}$, and the system will be located in a local minimum.
On the computational side, differential equations (35)-(36) are separable with respect to positions and momenta, and thus amenable to the numerically stable, symplectic Euler integration scheme, which can be efficiently implemented on GPUs. Moreover, the chaotic nature of the SBM equations imply sensitivity to initial conditions, and because each run is independent, many starting points can be evolved simultaneously, further increasing the efficiency and potential for massive parallelization.